OpenStudy (anonymous):

Simplify this expression : (x^27*y^62)/(x^13*y^30)

OpenStudy (anonymous):

Correct place to post? Complete solution: \[ \frac {x^{27} y^{62}}{x^{13} y^{30}} = \frac {x^{27}}{x^{13}} \cdot \frac {y^{62}}{y^{30}} = x^{27-13} y^{62-30} = \boxed {x^{14} y^{32}}. \]